# The Fate of Entanglement

### v2

@@ Abstract
Quantum entanglement is a fundamentally non-local correlation between particles. In its simplest realisation, a measurement on one particle is affected by a prior measurement on its partner, irrespective of their separation. For multiple particles, purely collective types of entanglement exist but their detection, even theoretically, remains an outstanding open question. Here, we show that all forms of multi-party entanglement entirely disappear during the typical evolution of a system as it heats up, evolves in time, or as its parts become separated. These results follow from the nature of the entanglement-free continent in the space of physical states, and hold in great generality. We illustrate these phenomena with a frustrated molecular quantum magnet in and out of equilibrium. In contrast, if the particles are fermions, such as electrons, another notion of entanglement exists that precludes entanglement-free regions, and thus protects quantum correlations. These findings provide fundamental knowledge about the structure of entanglement in quantum matter and architectures, paving the way for its manipulation.
@@
Introduction.— Entanglement is a non-local quantum correlation between two or more particles that makes it possible for a measurement on a subset of particles to affect subsequent measurements on the other particles. The effect of the initial measurement is instantaneous, even if the particles are distant. Entanglement not only constitutes a fundamental property of nature, but it is also a resource to perform tasks that would prove impossible without it such as teleportation [1], or more broadly quantum computation [2]. This has driven the community to devise methods to detect and quantify entanglement [3]. Unfortunately, it it not known how to determine with certainty whether a general system is entangled, except in very simple situations such as with 2 qubits. One can better grasp the complexity of the task by observing that entanglement can exist between more than two parties. In fact, some systems possess 3-party entanglement but no 2-party entanglement of any sort. In this work, we find criteria for such collective non-local quantum correlations in physical systems under very general conditions. For instance, at what temperatures can multi-party entanglement of a given kind exist? We begin by explaining important properties about the space of physical states, and how these determine the fate of entanglement under the evolution of a system with temperature, time or separation. We then illustrate these results with a simple yet generic model: the anti-ferromagnetic Ising model on an icosahedral molecule. Finally, we discuss the fate of entanglement for fermionic systems, where the parity superselection rule greatly modifies the geometry of the space of states and the structure of entanglement.
@@
Separable continent.— We investigate multi-party entanglement of states with $m$ subsystems, as illustrated in Fig. 1a. We shall argue that the end point of the evolution typically corresponds to an un-entangled state, which is called separable. The simplest separable state for a system of $m$ parties is a product,
@@
where $\rho_{j}$ is a physical density matrix for subsystem $j$. The set of separable sates is convex [3], namely states composed of a mixture of product states, $\rho_{\rm sep}=\sum_{k}p_{k}\rho_{\rm prod}^{(k)}$ with $p_{k}\geqslant 0$, are also un-entangled. In the space of all physical states, separable ones thus form a separable continent surrounded by an ocean of entangled states, see Fig. 1b. Highly-entangled states live in the deep-blue waters, whereas in the center of the separable continent lies the maximally-mixed state, i.e., the most classical state obtained in the limit of infinite temperature.
@@
We describe the system’s state by $\rho(s)$, where $s$ parametrizes the evolution; the final state is $\rho_{f}\equiv\rho(\infty)$. We shall argue that in situations of physical relevance the final state is generically separable. Moreover, $\rho_{f}$ typically lies in the interior of the separable continent, not on its frontier. We thus arrive at our main conclusion: the system irreversibly looses all forms of multi-party entanglement beyond some stage in the evolution, as shown in Fig. 1b. Moreover, before reaching land, the system navigates shallow waters, leading to a rapid decimation of entanglement. During this stage one has an effectively separable state. This constitutes a point of significant importance in real-world applications since determining when a state reaches a very low degree of entanglement is easier than showing exact separability. Nevertheless, we can exploit a mathematical result that tells us when a state is in the interior of the separable continent [4, 5, 6, 7]. In the simplest version, the theorem states that a product state (1) lies in the interior if it has full rank, i.e., none of its eigenvalues vanishes. This is coherent with what is known for pure product states: these have minimal rank (a single non-zero eigenvalue), and indeed live on the frontier of the continent as arbitrarily weak perturbations can make them entangled. Furthermore, the argument gives us the radius of a ball in the space of states that lies entirely on the separable continent; this is represented by the dashed line in Fig. 1b. The radius of the ball is proportional to the smallest eigenvalue of the state [7], $R_{m}=2^{1-m/2}\lambda_{\rm min}$. Here, we use the standard notion of distance between two quantum states given by the Frobenius norm: $d(\rho,\rho^{\prime})=\sqrt{\operatorname{Tr}(\rho-\rho^{\prime})^{2}}$. A more general version of the result holds for a mixture of product states with a least one having full rank [7].
@@
Figure 1: Evolution of multi-party entanglement. a) We consider a general state of $m$ subsystems, here illustrated for $m=4$, which can be in contact with an environment. b) The state is described by a density matrix $\rho(s)$ that evolves according to a parameter $s$ such as temperature, time or separation. The blue region represents the sea of entangled states, where deep-blue regions are more entangled than light-blue ones. The orange disk is the separable continent.
@@
Temperature.— In the first application, we take the parameter $s$ to represent the temperature $T$. Naturally, a very large temperature destroys entanglement, yielding a maximally uncorrelated state. The infinite-temperature end point is a full-rank product state, $\rho_{f}=D^{-1}\mathbb{I}$, where $\mathbb{I}$ represents the identity matrix in the $D$-dimensional space of the system. This state is located at the center of the separable continent. Hence, there exists a threshold temperature $T_{m}$ at which we can rigorously conclude that all forms of entanglement between the $m$ subsystems disappear. The minimal eigenvalue of $\rho_{f}$ being $1/D$, one can readily obtain a temperature above which entanglement is lost. However, in the example below we shall see that the system rapidly enters shallow waters, and becomes effectively separable at much lower temperatures. Sudden death of bipartite entanglement at finite temperature has been observed in different physical systems [8, 9, 10, 11, 12].
@@
Time.— In a dynamical situation, the parameter $s$ is the time $t$, and the final state is the state at $t=\infty$. If $\rho_{f}$ is on the separable continent, one can conclude that there is a sudden-death time $t^{*}_{m}$ at which all $m$-party entanglement is lost. As the structure of the final state depends on the type of dynamical evolution under study, let us describe an important case in more detail: a quantum quench. One prepares a closed system to be in an eigenstate of a given Hamiltonian $H_{I}$, and at some time the Hamiltonian is abruptly changed to a different one, $H$, resulting in non-trivial time evolution. We then study the state of a subset $A$ of the entire system: the $m$-party state $\rho(t)$ is obtained by partially tracing over the unobserved part, $B$. When $A$ is small compared to its complement $B$, the state is expected to effectively thermalize at large times. As discussed above, temperature tends to destroy entanglement, which leads to the expectation that numerous quench protocols will land on the continent.
@@
We can argue in full generality that the entanglement dynamics after a quench from a pure product state typically follows a rise-and-fall behavior. For early times, we consider the evolution of the system with parameter $s=t^{-1}$. The “final” state of this evolution is the initial state of the quench, namely a pure product state with rank one, i.e., not full rank. As we previously discussed, such states lie on the boundary, or the shore, of the separable continent. Therefore, there is no entanglement sudden death in this evolution. Looking back at the quench protocol with time as the parameter, this implies that entanglement is generated at $t=0^{+}$ after the quench. As discussed above, at late times the state typically lands on the continent, or in very shallow waters at a finite time, yielding an entanglement sudden death or strong suppression. These early- and late-time behaviors result in the rise-and-fall dynamics of entanglement. Such behavior has been observed for simple entanglement measures in numerous systems [13, 14, 15, 16, 17, 18, 19].
@@
Space.— We now study the fate of entanglement as a function of the separation between the $m$ subsystems by scaling their separations by $\lambda$, which parametrizes the evolution. At large $\lambda$, in a local physical system the state will factorise into a product form, where the description of each subsystem becomes independent, since all correlations become suppressed at large separations. If the asymptotic product state is of full rank, there exists a critical scale beyond which all entanglement vanishes. An important application is to study the entanglement between $m$ subsystems embedded in a larger system as these are chosen to be progressively further apart. The state at infinite $\lambda$ satisfies the product form, Eq. (1), where $\rho_{j}$ is the reduced density matrix of subsystem $j$ obtained by tracing out the complementary degrees of freedom. Let us take the entire system to be in equilibrium at a temperature $T~{}\!\!\geqslant~{}\!\!0$. At large separation, the state of subsystem $A_{1}$ is given by $\rho_{1}=\operatorname{Tr}_{A_{2}...A_{m}B}\sum_{n}p_{n}|E_{n}\rangle\langle E_{n}|$, where we have traced over the environment $B$, and the $(m-1)$ subsystems $A_{j}$. The sum runs over all eigenstates of the Hamiltonian. The thermal Boltzmann probabilities are $p_{n}=e^{-E_{n}/T}/\mathcal{Z}$. We now need to determine whether $\rho_{1}$ has full rank. Let us first examine the restriction of the eigenstates of the Hamiltonian to $A_{1}$, $\rho_{1}^{(n)}=\operatorname{Tr}_{A_{2}...A_{m}B}|E_{n}\rangle\langle E_{n}|$. From the point of view of $A_{1}$, the degrees of freedom in $B$ act as a bath, thus introducing statistical randomness in $\rho_{1}^{(n)}$. Since we take $A_{1}$ to be sufficiently small compared to the complement, the bath generically has sufficient resources to induce statistical fluctuations that span the entire Hilbert space of $A_{1}$. Such ergodicity implies that $\rho_{1}^{(n)}$ will be of full rank. This indeed occurs for the majority of eigenstates as they live near the middle of the energy spectrum. The eigenstate thermalization hypothesis [20, 21, 22] then states that the reduced density matrix on $A_{1}$ is approximately thermal, with the temperature determined by the energy $E_{n}$. The approximately thermal density matrix, obtained by restricting the Hamiltonian to $A_{1}$, then has full rank. Our statement, which we call the full-rank hypothesis (FRH), is more general: all eigenstates of a generic local Hamiltonian have a full-rank reduced density matrix associated with a sufficiently small subregion. In practice, the subregion should not exceed half the system.
@@
Now, the reduced density matrix of subsystem 1 is the convex sum $\rho_{1}=\sum_{n}p_{n}\rho_{1}^{(n)}$. The FRH implies that $\rho_{1}$ inherits full rank from the states $\rho_{1}^{(n)}$. (Since the sum of a full-rank matrix and an arbitrary one also has full rank, we actually only need to know that at least one $\rho_{1}^{(n)}$ is of full rank.) As the same holds for the other subsystems, we conclude that the $\lambda=\infty$ product has full rank, and thus lies on the separable continent. Therefore, there exists a scale beyond which all entanglement disappears. The sudden death occurs irrespective of the number of subsystems or the precise nature of the state; it even holds at a quantum critical phase transition where quantum fluctuations proliferate to all scales. This result shows that, although local observables can possess slowly decaying algebraic correlations, entanglement decays drastically faster, to the point of having a finite range. Our conclusion encompasses and generalises numerous examples of bipartite entanglement sudden death at finite separation [23, 24, 25, 26, 27, 28, 29], which are very specific cases of the general multi-party phenomenon.

### v4

@@ Abstract
Quantum entanglement manifests itself in non-local correlations between the constituents of a system. In its simplest realization, a measurement on one subsystem is affected by a prior measurement on its partner, irrespective of their separation. For multiple parties, purely collective types of entanglement exist but their detection, even theoretically, remains an outstanding open question. Here, we argue that all forms of multipartite entanglement entirely disappear during the typical evolution of a physical state as it heats up, evolves in time in a large family of dynamical protocols, or as its parts become separated. We focus on the generic case where the system interacts with an environment. These results mainly follow from the geometry of the entanglement-free continent in the space of physical states, and hold in great generality. We illustrate these phenomena with a frustrated molecular quantum magnet in and out of equilibrium, and a quantum spin chain. In contrast, if the particles are fermions, such as electrons, another notion of entanglement exists that protects bipartite quantum correlations. However, genuinely collective fermionic entanglement disappears during typical evolution, thus sharing the same fate as in bosonic systems. These findings provide fundamental knowledge about the structure of entanglement in quantum matter and architectures, paving the way for its manipulation.
I Introduction
@@
Entanglement is a fundamental property of quantum mechanics which can manifest itself in non-local quantum correlation between two or more systems. In particular, it makes it possible for a measurement on a subset of the whole system to affect subsequent measurements on the other parties. The effect of the initial measurement is instantaneous, even if the subsystems are distant. Entanglement not only constitutes a fundamental property of nature, but it is also a resource to perform tasks that would prove impossible without it such as teleportation [1, 2], or more broadly quantum computation [3, 4, 5]. This has driven the community to devise methods to detect and quantify entanglement [6, 7, 8]. Unfortunately, it it not known how to determine with certainty whether a general system is entangled, except in very simple situations such as with 2 qubits. One can better grasp the complexity of the task by observing that entanglement can exist between more than two parties. In fact, some systems possess 3-party entanglement but no 2-party entanglement of any sort, such as certain mixed states constructed from unextendible product bases [9].
In this work, we identify conditions for the existence of multipartite entanglement in quantum many-body systems such as spin models, the fermion Hubbard model, local quantum circuits, etc. We allow the state to be interacting with an environment: this can take the form of a thermal bath, or the neighboring spins/fermions when one considers a subregion. For instance, we answer the question: at what separations can multipartite entanglement of a given kind exist in a quantum many-body system/material? We begin by explaining important properties about the space of physical states, and how these determine the fate of entanglement under the evolution of a system with temperature, time or separation. We then illustrate these results with a simple yet generic model: the frustrated anti-ferromagnetic Ising model on an icosahedral molecule. Finally, we discuss the fate of entanglement for fermionic systems, where the fermion parity superselection rule modifies the geometry of the space of states and the structure of entanglement.
II Separable set
@@
We investigate multipartite entanglement of states with $m$ subsystems, as illustrated in Fig. 1a. We shall argue that in numerous physical situations, the end point of the evolution typically corresponds to an un-entangled state, which is called separable. The simplest separable state for a system of $m$ parties is a product,
@@
where $\rho_{j}$ is a physical density matrix for subsystem $j$. The set of separable sates is convex [6], namely states composed of a mixture of product states, $\rho_{\rm sep}=\sum_{k}p_{k}\rho_{\rm prod}^{(k)}$ with $p_{k}\geqslant 0$ and $\sum_{k}p_{k}=1$, are also un-entangled. In the space of all physical states, separable ones thus form a “continent” surrounded by an “ocean” of entangled states, see Fig. 1b. We stress however that this representation does not account for the intricate geometry of the boundary, but it illustrates the relevant ingredients for our discussion, namely the fact that the set of separable states is convex and has a non-vanishing volume. In our schematic representation, highly-entangled states live in the deep-blue regions, whereas in the center of the separable continent lies the maximally-mixed state, i.e., the identity.
@@
We describe the system’s state by $\rho(s)$, where $s$ parametrizes the evolution; the final state is $\rho_{f}\equiv\rho(s_{f})$, and we typically (but not always) consider $s_{f}=\infty$. We shall argue that in numerous situations of physical relevance the final state is separable. Moreover, $\rho_{f}$ typically lies in the interior of the separable continent, not on its frontier. Under these conditions, we thus arrive at our main conclusion: in numerous physically relevant situations, the system irreversibly looses all forms of multipartite entanglement beyond some stage in the evolution, as shown in Fig. 1b. Moreover, before reaching land, the system navigates shallow waters, leading to a rapid decimation of entanglement. During this stage one has an effectively separable state. This constitutes a point of significant importance in real-world applications since determining when a state reaches a very low degree of entanglement is easier than showing exact separability. Indeed, numerical estimations of the degree of entanglement typically involve optimization procedures which correctly yield small results for low-entangled states, but do not necessarily perfectly detect exact separability. This phenomenon notably occurs for the geometric entanglement, discussed later in the text. Nevertheless, we can exploit a mathematical result that tells us when a state is in the interior of the separable continent [10, 11, 12, 13]. In the simplest version, the theorem states that a product state (1) lies in the interior if it has full rank, i.e., none of its eigenvalues vanishes. This is coherent with what is known for pure product states: these have minimal rank (a single non-zero eigenvalue), and indeed live on the frontier of the continent as arbitrarily weak perturbations can make them entangled. Furthermore, the theorem yields the radius of a ball in the space of states that lies entirely on the separable continent; this is represented by the dashed line in Fig. 1b. The radius of the ball is proportional to the smallest eigenvalue of the state [13], $R_{m}=2^{1-m/2}\lambda_{\rm min}$. Here, we use the standard notion of distance between two quantum states given by the Frobenius norm: $d(\rho,\rho^{\prime})=\sqrt{\operatorname{Tr}(\rho-\rho^{\prime})^{2}}$. A more general version of the result holds for a mixture of product states with a least one having full rank [13]. This implies that the boundary of the separable continent does not only contain pure product states, but also higher-rank and full-rank separable states whose product-state components are all rank-deficient.
II.1 Biseparable set and genuine multipartite entanglement
In multiparty states, there is genuine multipartite entanglement (GME) if the state is not biseparable, i.e., if it cannot be written as a convex combination of product states over bipartitions of the $m$ parties,
$\rho_{\rm bisep}=\sum_{k}p_{k}\ \rho_{\mathcal{I}_{k}}\otimes\rho_{\overline{\mathcal{I}}_{k}},$ (2)
where $\mathcal{I}_{k}\cup\overline{\mathcal{I}}_{k}$ are bipartitions labelled by $k$, $p_{k}\geqslant 0$, and $\sum_{k}p_{k}=1$. As an example, for $m=3$ parties, the possible bipartitions are $(1,23)$, $(12,3)$ and $(13,2)$. Biseparable states contain at most $(m-1)$-party entanglement, which is not genuinely multipartite. Similarly to separable states, biseparable ones form a convex set which includes the separable continent. Using the same arguments as above, we conclude that during typical generalized evolutions $\rho(s)$, the state irreversibly loses GME at a stage prior to the sudden death of entanglement, implying that non-GME and bipartite entanglement are more robust than GME.
@@ II.1 Biseparable set and genuine multipartite entanglement
Figure 1: Evolution of multipartite entanglement. a) We consider a general state of $m$ subsystems, here illustrated for $m=4$, which can be in contact with an environment. b) The state is described by a density matrix $\rho(s)$ that evolves according to a parameter $s$ such as temperature, time or separation. The blue region represents the “sea” of entangled states, where deep-blue regions are more entangled than light-blue ones. The orange disk is the separable continent.
II.2 Temperature
@@
As a warmup case that will be a reference point when we shall discuss thermalisation, we take the parameter $s$ to represent the temperature $T$, as would pertain to a Gibbs thermal state associated with a Hamiltonian $H$, $\rho(T)=\exp(-H/T)/\mathcal{Z}$, or a reduced density matrix thereof $\rho_{A}(T)=\operatorname{Tr}_{B}\rho(T)$. Naturally, a very large temperature destroys entanglement, yielding a maximally uncorrelated state. The infinite-temperature end point is a full-rank product state, $\rho_{f}=D^{-1}\mathbb{I}$, where $\mathbb{I}$ represents the identity matrix in the $D$-dimensional space of the system. This state is located at the center of the separable continent. Hence, there exists a threshold temperature $T_{m}$ at which one can rigorously conclude that all forms of entanglement between the $m$ subsystems disappear, in agreement with previous results for finite-dimensional systems [14]. At the threshold, the state at the boundary of the separable continent will generically have full-rank albeit being in a mixture of only rank-deficient product states. The minimal eigenvalue of $\rho_{f}$ being $1/D$, one can readily obtain an upper bound for the temperature above which entanglement is lost. However, the example below will show that physical many-body systems tend to loose entanglement much faster than this upper bound. Sudden death of bipartite entanglement at finite temperature has been observed in different physical systems [15, 16, 17, 18, 19]. Later, we will contrast the above behavior with the thermal fate of entanglement in fermion systems.
II.3 Time
@@
In a dynamical situation, the parameter $s$ is the time $t$, and we take the final state to be the state at some time $t_{f}$. In dynamical protocols where the system reaches a final state $\rho_{f}$ which lies inside the separable continent, one can conclude that there is a sudden-death time $t^{*}_{m}$ at which all $m$-party entanglement is lost until $t_{f}$. As the structure of the final state depends on the type of dynamical evolution under study, let us describe an important case in more detail: a quantum quench. One prepares a closed system to be in an eigenstate of a given Hamiltonian $H_{I}$, and at some time the Hamiltonian is abruptly changed to a different one, $H$, resulting in non-trivial time evolution. We then study the state of a subregion $A$ of the entire system: the $m$-party state $\rho(t)$ is obtained by partially tracing over the unobserved part, $B$. When $A$ is small compared to its complement $B$, the state is expected to effectively thermalize at large times [20, 21, 22]. As discussed above, temperature tends to destroy entanglement, which means that in numerous quench protocols the subregion final state $\rho(t_{f})$ will be separable. Although not all quenches will lead to a sudden death of multipartite entanglement, one can ask about the typicality of such a fate. For instance, if the initial condition has an energy expectation value close to the groundstate of $H$, instead of thermalisation, oscillatory behavior with persistent entanglement can occur. However, only a minority of eigenstates of a local Hamiltonian $H$ have an energy near the minimum (or maximum) of the spectrum: the bulk of the levels lies near the middle of the spectrum. As such, a typical initial condition will inject enough energy so that the evolution will scramble the quantum information encoded within $A$. In many cases one can be more precise: the state of $A$ at large times can is accurately approximated by an effective thermal Gibbs state, $\operatorname{Tr}_{B}(\exp(-H/T_{\rm eff}))/\mathcal{Z}$, When the effective temperature $T_{\rm eff}$ is not small (which is typical), an entanglement sudden death for the subregion follows. Below we shall rigorously verify the dynamical sudden death with numerous examples.
@@
To complement the discussion of late-time dynamics, we can argue in full generality that the multiparty entanglement dynamics after a quench from a pure product state typically follows a rise-and-fall behavior. For early times, we consider the evolution of the system with parameter $s=t^{-1}$. The “final” state of this evolution is the initial state of the quench, namely a pure product state with rank one, i.e., not full rank. As we previously discussed, such states lie on the boundary, or the shore, of the separable continent. Therefore, there is no entanglement sudden death in this evolution. Looking back at the quench protocol with time as the parameter, this implies that entanglement is generated at $t=0^{+}$ after the quench. As discussed above, in numerous situations, the state typically lands on the continent at late times, or in very shallow waters at a finite time, yielding an entanglement sudden death or strong suppression. These early- and late-time behaviors result in the rise-and-fall dynamics of entanglement, which have been observed for simple entanglement measures in various systems [23, 24, 25, 26, 27, 28, 29].
II.4 Space
@@
We now study the fate of entanglement as a function of the separation between the $m$ subsystems by scaling their separations by $\lambda$, which parametrizes the evolution. At large $\lambda$, in a local physical system the state will factorise into a product form, where the description of each subsystem becomes independent, since all correlations become suppressed at large separations. If the asymptotic product state is of full rank, there exists a critical scale beyond which all entanglement vanishes. As a central application, we consider a quantum many-body system from which we extract the state of a subregion whose $m$ parts, $A=A_{1}\cdots A_{m}$, become more separated as $\lambda$ grows, $\rho(\lambda)$. The state at infinite $\lambda$ satisfies the product form, Eq. (1), where $\rho_{j}$ is the reduced density matrix of subsystem $j$ obtained by tracing out the complementary degrees of freedom. Let us take the entire system to be in equilibrium at a temperature $T~{}\!\!\geqslant~{}\!\!0$, described by the Gibbs state $\rho_{AB}=\mathcal{Z}^{-1}\exp(-H/T)$. At large separation, the state of subsystem $A_{1}$ is given by $\rho_{1}=\operatorname{Tr}_{A_{2}...A_{m}B}\sum_{n}p_{n}|E_{n}\rangle\langle E_{n}|$, where we have traced over the environment $B$, and the $(m-1)$ subsystems $A_{j}$. The sum runs over all eigenstates of the Hamiltonian $H$. The thermal Boltzmann probabilities are $p_{n}=e^{-E_{n}/T}/\mathcal{Z}$. We now need to determine whether $\rho_{1}$ has full rank. Let us first examine the restriction of the eigenstates of the Hamiltonian to $A_{1}$, $\rho_{1}^{(n)}=\operatorname{Tr}_{A_{2}...A_{m}B}|E_{n}\rangle\langle E_{n}|$. From the point of view of $A_{1}$, the degrees of freedom in the complement act as a bath, thus introducing statistical randomness in $\rho_{1}^{(n)}$. Since we take $A_{1}$ to be sufficiently small compared to the complement and the interactions generic (not fine-tuned), the bath generically has sufficient resources to induce statistical fluctuations that span the entire Hilbert space of $A_{1}$. Such ergodicity implies that $\rho_{1}^{(n)}$ will be of full rank. For instance, it is readily seen to occur for the majority of eigenstates as they live near the middle of the energy spectrum, and the eigenstate thermalization hypothesis [30, 31, 32] then states that the reduced density matrix on $A_{1}$ is approximately thermal, with the temperature determined by the energy $E_{n}$. The approximately thermal density matrix, obtained by restricting the Hamiltonian to $A_{1}$, then has full rank. Our statement, which we call the full-rank hypothesis (FRH), is more general: all eigenstates of a generic local Hamiltonian have a full-rank reduced density matrix associated with a sufficiently small subregion. In practice, the subregion should not exceed half the system. To complement the physical argument given above, we shall give an example in a frustrated quantum magnet in the next section.
@@
Now, the reduced density matrix of subsystem 1 is the convex sum $\rho_{1}=\sum_{n}p_{n}\rho_{1}^{(n)}$. The FRH implies that $\rho_{1}$ inherits full rank from the states $\rho_{1}^{(n)}$. (Since the sum of a full-rank matrix and an arbitrary one also has full rank, we actually only need to know that at least one $\rho_{1}^{(n)}$ is of full rank.) As the same holds for the other subsystems, we conclude that the $\lambda=\infty$ product has full rank, and thus lies inside the separable continent (not on its frontier). Therefore, there exists a scale beyond which all entanglement disappears. The sudden death occurs irrespective of the number of subsystems or the precise nature of the state; it even holds at a quantum critical phase transition where quantum fluctuations proliferate to all scales. This result shows that, although local observables can possess slowly decaying algebraic correlations, entanglement decays drastically faster, to the point of having a finite range. This can be understood from the monogamy of entanglement: in a typical local Hamiltonian, the interactions favour entanglement among nearby degrees of freedom, which strongly limits the capacity to entangle with more distant sites. In analogy with charged impurities in a material, one could say that “entanglement is screened” by the environment of $A=A_{1}\cdots A_{m}$. Our conclusion encompasses and generalises numerous examples of bipartite entanglement sudden death at finite separation [33, 34, 35, 36, 37, 38, 39], which are very specific cases of the general multipartite phenomenon. Below we illustrate the rigorous finite range of multiparty entanglement in two models.
III Quantifying Entanglement
In order to illustrate the Fate of Entanglement (FoE) under various evolutions, we will evaluate powerful quantifiers of entanglement, including for genuine multiparty entanglement. First, we compute a measure of 2-party entanglement, focusing on a pair of adjacent spins, called the logarithmic negativity $\mathcal{E}$ [40, 41]. In the case of two spins, $\mathcal{E}=0$ implies that the reduced density matrix is separable [42], whereas for entangled states we have $\mathcal{E}>0$. Second, we study multiparty entanglement via the geometric entanglement $\mathcal{D}$ [6] defined as the smallest Hilbert-Schmidt distance between the state and the separable set,
$\mathcal{D}=\min_{\sigma\in\,{\rm SEP}}d(\rho,\sigma),$ (3)
where the minimization is over all $\rho_{\rm sep}$ living in the set of separable states of $m$-parties. Similarly, one can define the geometric genuine multiparty entanglement as
$\mathcal{D}_{b}=\min_{\sigma\in\,{\rm bSEP}}d(\rho,\sigma),$ (4)
where SEP and bSEP refer to the separable and biseparable sets, respectively. These are powerful quantities: $\mathcal{D}$ detect all forms of entanglement, while $\mathcal{D}_{b}$ detects all genuine multiparty entanglement. However, they can be difficult to evaluate due to the optimisation over a large parameter space. Nevertheless, an efficient method (due to Gilbert) exploiting the convexity of SEP and bSEP was shown to produce strong results. We will also employ a more direct global optimisation procedure, which when used in a “layered” fashion can produce better upper bounds on $\mathcal{D},\mathcal{D}_{b}$ in certain cases. The values we give below are the best of the two methods for the given state, and satisfy convergence criteria (see Appendix). Finally, we also employ a much simpler criterion $W$ which indicates the presence of genuine 3-spin entanglement by detecting a property that cannot hold for biseparable states (2[43]. It is defined from the matrix elements $\rho_{ij}$, $i,j=1,\dots,8,$ of a 3-spin density matrix [43]:
$W=|\rho_{23}|+|\rho_{25}|+|\rho_{35}|-\sqrt{\rho_{11}\rho_{44}}-\sqrt{\rho_{11}\rho_{66}}-\sqrt{\rho_{11}\rho_{77}}-\frac{1}{2}(\rho_{22}+\rho_{33}+\rho_{55}).$ (5)
$W>0$ indicates the presence of genuine 3-party entanglement, whereas for $W\leqslant 0$ a definitive conclusion cannot be made. We thus set $W=0$ in this case. As the above expression for the criterion is basis-dependent, we maximize its value over all possible local unitary transformations of the state, $(U_{1}\otimes U_{2}\otimes U_{3})\rho(U_{1}^{\dagger}\otimes U_{2}^{\dagger}\otimes U_{3}^{\dagger})$, where $U_{j}$ is a generic $2\times 2$ unitary matrix for the spin $j$. While it is not a proper entanglement measure, large values of $W$ generically correspond to stronger genuine 3-party entanglement. For instance, $W$ is maximal for Werner states, which are maximally-entangled 3-qubit states, and is minimal for the maximally mixed state (the identity).
In the cases considered below, we obtain values of $\mathcal{D}_{(b)}$ on the order of $10^{-6}$ beyond a certain point in the evolution, which implies that the corresponding type of entanglement is so small that it is of no practical relevance. One could stop there and move on. However, we want to make rigorous statements regarding the sudden death in order to demonstrate the general results in the strongest terms. We will thus employ an approach that can rigorously certify whether a density matrix is (bi)separable: the general trace criterion (GTC) [44]. The idea is to use filtering and subtractions to bring the original state into an ellipsoid of (bi)separable states around a reference state; the ellipsoid is strictly larger than the ball referred to above. The numerical optimisation used is very similar to the one for $\mathcal{D}_{(b)}$.
IV Icosahedral molecule
We first illustrate the above results with a simple quantum system: the anti-ferromagnetic Ising model on the 12-spin icosahedron, see Fig. 2a, with Hamiltonian
$H=J\sum_{{\rm bonds}\,\langle i,j\rangle}\sigma_{i}^{x}\sigma_{j}^{x}-h\sum_{{\rm sites}\,i}\sigma_{i}^{z}$ (6)